Please help.

Solve by the elimination method.
3x + 4y=3
6x + 8y=6
What is the solution of the system?
(Type an ordered pair. Type an integer or a fraction. Type N if there is no solution. Type I if there are infinitely more solutions.)



Answer by checkley77(12844)

3x + 4y=3 multiply this equation by -2 & add.
6x + 8y=6
-6x-8y=-6
---------------
0x+y=0
N no solution.